# LightMDETR: A Lightweight Approach for Low-Cost Open-Vocabulary Object Detection Training

### v1

@@ Abstract
Object detection in computer vision traditionally involves identifying objects in images. By integrating textual descriptions, we enhance this process, providing better context and accuracy. The MDETR model significantly advances this by combining image and text data for more versatile object detection and classification. However, MDETR’s complexity and high computational demands hinder its practical use. In this paper, we introduce Lightweight MDETR (LightMDETR), an optimized MDETR variant designed for improved computational efficiency while maintaining robust multimodal capabilities. Our approach involves freezing the MDETR backbone and training a sole component, the Deep Fusion Encoder (DFE), to represent image and text modalities. A learnable context vector enables the DFE to switch between these modalities. Evaluation on datasets like RefCOCO, RefCOCO+, and RefCOCOg demonstrates that LightMDETR achieves superior precision and accuracy.
@@ I Introduction
Object detection is a key task in computer vision, involving identifying and localizing objects within images. Traditionally, closed-vocabulary models have been used, where models are trained to recognize a fixed set of object categories. Notable approaches like Faster R-CNN [1], YOLO [2], and SSD [3] have demonstrated high effectiveness but are limited in generalizing beyond predefined categories.
Open-vocabulary models offer new possibilities by using large, diverse datasets and unsupervised learning to detect objects not in the original training categories. Models like CLIP [4], ALIGN [5], and Florence [6] use transformer architectures to learn combined text and image representations, allowing for more flexible and context-sensitive object recognition.
In this context, open-vocabulary models like Multimodal Detr (MDETR) integrates image and text sequences, utilizing transformers to align visual and textual data for more versatile and accurate detection and classification. Similarly, Grounded Language-Image Pre-training (GLIP) [7] enhances detection performance by leveraging large-scale vision-language data. Models like RegionCLIP [8] and OWL-ViT [9] further advance multimodal learning by improving visual-textual alignment. These advancements highlight the potential of open-vocabulary models to revolutionize object detection, making it more adaptable to diverse and dynamic environments.
However, open-vocabulary models face significant challenges due to their complexity and high computational requirements, limiting their practical deployment in real-time applications or on resource-constrained devices.
To address these issues, we propose Lightweight MDETR (LightMDETR), an optimized version of the MDETR architecture. LightMDETR enhances computational efficiency while maintaining robust multimodal integration for object detection. Key optimizations include freezing the backbone components of the pre-trained model and introducing a ”Deep Fusion Encoder” (DFE) to represent text and image modalities using shared parameters. By integrating learnable ”context” parameters into each embedding, the DFE can effectively encode both modalities. This approach significantly reduces the number of parameters to tune while preserving the performance of the baseline MDETR model.
@@
  • We introduce a lightweight method for open-vocabulary object detection that significantly reduces the number of parameters to tune, making training more cost-effective.
  • We apply this approach to the MDETR architecture with two variants: LightMDETR, which trains only the ”Deep Fusion Encoder” (DFE), and LightMDETR-CF, which extends LightMDETR with a cross-fusion layer between text and image modalities to enhance representation capability.
  • We train only the DFE, freezing all pre-trained specialized backbone encoders for images and text. We achieve this by incorporating ”context” parameters into the DFE, allowing it to switch between image and text modalities effectively.
@@ II Related Work
II-A Evolution of Object Detection: From Closed-Vocabulary to Open-Vocabulary Models
Object detection, a central task in computer vision, can be broadly categorized into closed-vocabulary and open-vocabulary methods.
Closed-vocabulary object detection focuses on detecting and classifying objects from a predefined, fixed set of categories. Notable models in this category include Faster R-CNN [1], which introduced the Region Proposal Network (RPN) to generate object proposals directly, making the process more efficient. YOLO [2] brought a novel approach by framing detection as a single regression problem, enabling real-time object detection by predicting bounding boxes and class probabilities in a single pass. SSD [3] improved detection accuracy, especially for small objects, by introducing multi-scale feature maps for predictions. RetinaNet [10] addressed the challenge of class imbalance in object detection through its innovative Focal Loss, which down-weights the loss assigned to well-classified examples, enhancing the performance on hard-to-detect objects. Mask R-CNN [11] extended the Faster R-CNN framework by adding a parallel branch for predicting object masks, enabling instance segmentation in addition to object detection. While these models are highly effective within their defined scope, their limitation lies in their inability to generalize beyond the fixed set of classes, making them less adaptable in dynamic environments.
@@ II-A Evolution of Object Detection: From Closed-Vocabulary to Open-Vocabulary Models
Open-vocabulary object detection aims to transcend the constraints of fixed class categories by leveraging large-scale pretrained models that understand both text and image modalities. ViLD [12] uses CLIP’s embeddings to enable zero-shot object detection by matching image regions with any text description, a significant leap in flexibility and generalization. GLIP [7] integrates grounding into the pretraining process, allowing the model to learn the alignment between language and image regions more effectively. MDETR [13] combines textual cues dynamically with image features to improve detection in context-rich environments. Contextual Object Detection [14] emphasizes the role of surrounding context in improving detection accuracy, pushing the boundaries of what objects can be recognized by understanding relationships within the scene. Despite their superior generalization capabilities, these open-vocabulary methods are resource-intensive, requiring substantial computational power and large-scale datasets for training, primarily due to the reliance on extensive pretrained models for text and image encoding. Nonetheless, they represent a significant advancement in object detection, offering the ability to detect a vast range of objects, including those unseen during training.
@@
To tackle the challenges associated with the extensive training required for open-vocabulary object detection, we propose a new method that significantly reduces training demands while maintaining performance. Our approach can be seamlessly integrated into any existing open-vocabulary object detection model, ensuring more efficient training without compromising the model’s effectiveness. To demonstrate the effectiveness of our method, we validate it using the MDETR architecture.
II-B MDETR: Modulated Detection for End-to-End Multi-Modal Understanding
@@
MDETR advances object detection by integrating both visual and textual information into a unified framework. Unlike traditional object detection methods that classify objects into fixed categories, MDETR focuses on associating detected objects with spans of text tokens. The model uses ResNet [15] for visual feature extraction and RoBERTa [16] for textual feature extraction (ref. Figure 1). For training, MDETR employs two key loss functions to align image and text data. The soft token prediction loss ($\mathcal{L}_{soft\_token}$) guides the model to predict a uniform distribution over the tokens in the text that correspond to each detected object, rather than predicting discrete class labels.
Figure 1: MDETR Architecture: Visual features are extracted using ResNet, while textual features are extracted using RoBERTa.
@@
III Proposed Method
Existing open-vocabulary object detection methods often rely on large, specialized pre-trained models to separately encode images and text—such as using ResNet for image encoding and RoBERTa for text encoding, as shown in Figure 1. Training these large models can be prohibitively expensive, which limits the practicality of open-vocabulary methods compared to their closed-vocabulary counterparts. To address this issue and reduce training costs while preserving performance, we propose a lightweight training approach. This method allows for cost-effective training of any open-vocabulary model without sacrificing performance. We validate our approach by applying it to MDETR, resulting in a new variant called LightWeight MDETR (LightMDETR).
LightMDETR addresses the challenge of high training costs by freezing pre-trained feature extraction models, such as ResNet and RoBERTa, and introducing a single, lightweight component we developed called the ”Deep Fusion Encoder” (DFE) $f$. The DFE integrates embeddings from both image and text encoders, streamlining the process while maintaining performance, as illustrated in Figure 2.
Figure 2: LightMDETR architecture: ResNet and RoBERTa are frozen for feature extraction, with only the Deep Fusion Encoder (DFE) $f$ being trained. A learnable context vector $c_{l}$ where $l\in\{\text{image},\text{text}\}$, is incorporated into the ResNet and RoBERTa embeddings, enabling the DFE to share parameters while switching between modalities. The rest of the architecture, including the ”Transformer” module from DETR, remains unchanged from the baseline MDETR.
To allow the DFE to encode inputs from different models, such as ResNet and RoBERTa, using the same parameters, a learnable context vector $c_{l}$, where $l\in\{\text{image},\text{text}\}$, is fused with the embeddings. This approach enables the DFE to switch between image and text modalities while maintaining consistent parameter usage.
[…]
$O_{f}$ $=f(O\otimes c_{\text{image}})$ (8)
$T_{f}$ $=f(T\otimes c_{\text{text}})$
where $\otimes$ denotes the fusion operation (e.g., addition, multiplication, concatenation, or cross-attention), and $c_{\text{image}}$ and $c_{\text{text}}$ are randomly initialized vector parameters that match the dimensions of $O_{f}$ and $T_{f}$, respectively. These vectors are learned as weights during the backpropagation process.
After the DFE processes the representations, the resulting features $O$ and $T$ are concatenated and fed into the DETR [17] ”Transformer” (as shown in Figure 2). This enables the detection of relevant objects in the image, conditioned by the accompanying text.
To maintain a lightweight architecture, the DFE consists of two key components:
  • A fusion block: which combines the context vector with the embeddings from the pretrained encoders.
  • A single transformer layer: which refines and enhances the representation of the fused input, integrating both the embedding and context vector.
The loss function remains unchanged from MDETR (see Equation 4), allowing the learnable context vectors to be effectively adapted to the target task. This is achieved by directly contributing to the minimization of the loss function.
In MDETR, image and text features are encoded separately and only fused at the end before being input to the DETR. However, as noted in [18], effective phrase grounding in vision-language models requires early fusion of visual and textual features. To address this, we enhance LightMDETR by incorporating deep fusion of image and text features prior to DFE encoding, as illustrated in Figure 3. This modified approach is termed ”Cross-Fusion LightMDETR” (LightMDETR-CF).
@@ III Proposed Method
Figure 3: Architecture of LightMDETR-CF: LightMDETR-CF extends LightMDETR 2 by introducing a cross-fusion layer prior to DFE encoding, thereby enhancing the model’s representation capabilities.
LightMDETR-CF incorporates three key components into LightMDETR: a cross-fusion layer featuring Multi-Head Attention (MHA) [19], and two additional transformer layers that refine the MHA outputs before they are fed into the DFE.
@@
The MHA mechanism takes as input the outputs of the ResNet and RoBERTa encoders, denoted as $O$ and $T$, respectively. The transformations can then be expressed as follows:
@@
$T^{(v)}$ $=TW^{(v,T)},$ $O_{\text{CF}}$ $=\text{SoftMax}(\text{Attn})\cdot T^{(v)}\cdot W^{(\text{out},O)},$
$O^{(v)}$ $=OW^{(v,O)},$ $T_{\text{CF}}$ $=\text{SoftMax}(\text{Attn}^{\top})\cdot O^{(v)}\cdot W^{(\text{out},T)},$
@@
After applying the cross-fusion mechanism with the Multi-Head Attention approach, a projection is performed on $O_{\text{CF}}$ and $T_{\text{CF}}$ using two distinct Transformer layers:
[…]
The resulting $O_{TL}$ and $T_{TL}$ are then fed into the DFE, following a similar process as in LightMDETR, as described by:
$O_{f}$ $=f(O_{TL}\otimes c_{\text{image}}),$ (11)
$T_{f}$ $=f(T_{TL}\otimes c_{\text{text}}).$
These two approaches, LightMDETR and its variant LightMDETR-CF, provide an efficient strategy for training open-vocabulary object detection models. They significantly reduce training costs while maintaining high performance.
@@ IV-A Pre-training
For both the LightMDETR and LightMDETR-CF models, we use a frozen, pretrained RoBERTa-base [16] as the text encoder, which has 12 transformer layers, each with a hidden dimension of 768 and 12 attention heads, totaling 125M parameters. The visual backbone is a frozen, pretrained ResNet-101 [15], which has 44M parameters. By freezing both encoders, we reduce the trainable parameters in the backbone from 169 million in the original MDETR to zero in LightMDETR and LightMDETR-CF. Instead, the only trainable component in these models is the DFE (ref. Figure 2 and 3), which consists of a single transformer layer with 4 attention heads, amounting to 787,968 parameters to train.
For the fusion operation in the DFE, as described in Equation 11, we use an addition method. The updated equations are:
$O_{f}$ $=f(O+c_{\text{image}})$ (12)
$T_{f}$ $=f(T+c_{\text{text}})$
In these equations, $c_{\text{image}}$ and $c_{\text{text}}$ are initialized using a normal distribution and have the same dimension as $O$ and $T$, which is 256.
All models undergo pre-training for 40 epochs with a substantial effective batch size of 64.
@@ IV-B Dowstream Tasks
Models (LightMDETR and LightMDETR-CF) are evaluated in referring expression comprehension on RefCOCO [23], RefCOCO+ [24] and RefCOCOg [25] (ref. Table I. Referring expression comprehension involves the ability to accurately identify and localize objects in an image based on detailed natural language descriptions. This task requires interpreting the referring expressions—such as ”the blue umbrella next to the park bench”—to extract specific attributes and spatial relationships. The goal is to map these textual descriptions to the corresponding objects in the image, ensuring precise object detection and localization by integrating both natural language processing and computer vision techniques.
Datasets Images Instances Annotations Categories Image Size Vocab. Size
RefCOCO 3,000 7,596 21,586 71 230 - 640 3,525
RefCOCO+ 3,000 7,578 21,373 71 230 - 640 4,387
RefCOCOg 3,900 7,596 14,498 78 277 - 640 5,050
TABLE I: Datasets for validating referring expression comprehension.
Similar to MDETR, LightMDETR, and LightMDETR-CF are trained to predict bounding boxes for all objects mentioned in referring expressions, such as ”blue umbrella” and ”park bench.” However, the task in referring expression comprehension is to return a single bounding box corresponding to the object described by the entire expression. To address this, we fine-tune the model on a task-specific dataset for 5 epochs.

### v3

# A Lightweight Modular Framework for Low-Cost Open-Vocabulary Object Detection Training

@@ Abstract
Object detection is a fundamental challenge in computer vision, centered on recognizing objects within images, with diverse applications in areas like image analysis, robotics, and autonomous vehicles. Although existing methods have achieved great success, they are often constrained by a fixed vocabulary of objects. To overcome this limitation, approaches like MDETR have redefined object detection by incorporating region-level vision-language pre-training, enabling open-vocabulary object detectors. However, these methods are computationally heavy due to the simultaneous training of large models for both vision and language representations. To address this, we introduce a lightweight framework that significantly reduces the number of parameters while preserving, or even improving, performance. Our solution is applied to MDETR, resulting in the development of Lightweight MDETR (LightMDETR), an optimized version of MDETR designed to enhance computational efficiency without sacrificing accuracy. The core of our approach involves freezing the MDETR backbone and training only the Universal Projection module (UP), which bridges vision and language representations. A learnable modality token parameter allows the UP to seamlessly switch between modalities. Evaluations on tasks like phrase grounding, referring expression comprehension, and segmentation show that LightMDETR not only reduces computational costs but also outperforms several state-of-the-art methods in terms of accuracy. Our code implementation is available on our GitHub repository: LightMDETR github.
@@ I Introduction
Object detection is a crucial task in computer vision, focused on identifying and localizing objects within images. Leading methods like Faster R-CNN [1], YOLO [2], and SSD [3] have shown great success in this domain. However, these approaches are constrained by a fixed set of object categories (e.g., 20 categories in the PASCAL VOC [4] dataset). Once trained, these detectors can only recognize the predefined categories, limiting their flexibility and applicability in more open and dynamic scenarios.
Recent works [5, 6, 7, 8] have leveraged popular vision-language models for open-vocabulary detection by distilling vocabulary knowledge from language encoders. However, these distillation-based approaches face significant limitations due to the scarcity of diverse training data.
Inspired by the success of methods [9, 10, 11] that learn image-level visual representations from large-scale raw image-text pairs, achieving semantically rich projection spaces for easy transfer to downstream tasks (such as zero-shot image classification and text-image retrieval), several approaches [12, 13, 7, 14, 15] have extended this to open-vocabulary object detection, aiming for fine-grained image understanding with object-level visual representations. Despite their progress, these methods still suffer from heavy computational overhead, as they rely on large models to encode both visual and language data.
To tackle this challenge, we introduce a lightweight modular framework that can be seamlessly incorporated into any open-vocabulary object detection system, reducing training costs by minimizing the number of tunable parameters, while preserving or even boosting the baseline object detector’s performance. The key innovations of this approach include freezing the backbone of pre-trained models and introducing a ”Universal Projection” (UP) module that shares parameters to represent both visual and language data. To ensure the UP can handle data from different distributions (visual and language) using the same parameters, we incorporate a learnable ”modality token” that enables the UP to switch between the two modalities effectively. We integrate this framework into MDETR, resulting in Lightweight MDETR (LightMDETR), and validate its effectiveness on tasks such as phrase grounding, referring expression comprehension, and segmentation.
@@
  • We introduce a lightweight method for open-vocabulary object detection systems that significantly reduces the number of parameters to tune, making training more cost-effective.
  • We apply this approach to the MDETR architecture with two variants: LightMDETR, which trains only the ”Universal Projection” (UP) module, and LightMDETR-CF, which extends LightMDETR with a cross-fusion layer between text and image modalities to enhance representation capability.
  • We train only the UP, freezing all pre-trained specialized backbone models for images and text. We achieve this by incorporating ”modality token” parameter into the UP, allowing it to switch between image and text modalities effectively.
@@ II Related Work
II-A Traditional Object Detection
Traditional object detection focuses on detecting and classifying objects from a predefined fixed set of categories. Notable models in this category include Faster R-CNN [1], which introduced the Region Proposal Network (RPN) to generate object proposals directly, making the process more efficient. YOLO [2] brought a novel approach by framing detection as a single regression problem, enabling real-time object detection by predicting bounding boxes and class probabilities in a single pass. SSD [3] improved detection accuracy, especially for small objects, by introducing multi-scale feature maps for predictions. RetinaNet [16] addressed the challenge of class imbalance in object detection through its innovative Focal Loss, which down-weights the loss assigned to well-classified examples, enhancing the performance on hard-to-detect objects. Mask R-CNN [17] extended the Faster R-CNN framework by adding a parallel branch for predicting object masks, enabling instance segmentation in addition to object detection. While these models are highly effective within their defined scope, their limitation lies in their inability to generalize beyond the fixed set of categories, making them less adaptable in dynamic environments.
@@ II-B Open-vocabulary Object Detection (OVD)
OVD aims to transcend the constraints of fixed class categories by leveraging large-scale pre-trained models that understand both text and image modalities. ViLD [7] uses CLIP’s embeddings to enable zero-shot object detection by matching image regions with any text description, a significant leap in flexibility and generalization. GLIP [18] integrates grounding into the pre-training process, allowing the model to learn the alignment between language and image regions more effectively. MDETR [12] combines textual cues dynamically with image features to improve detection in context-rich environments. Contextual Object Detection [19] emphasizes the role of surrounding context in improving detection accuracy, pushing the boundaries of what objects can be recognized by understanding relationships within the scene. Despite their superior generalization capabilities, these open-vocabulary methods are resource-intensive, requiring substantial computational power and large-scale datasets for training, primarily due to the reliance on extensive pre-trained models for text and image encoding. Nonetheless, they represent a significant advancement in object detection, offering the ability to detect a vast range of objects, including those unseen during training.
@@
To tackle the challenges associated with the extensive training required for open-vocabulary object detection methods, we propose a new method that significantly reduces training demands while maintaining performance. Our approach can be seamlessly integrated into any existing open-vocabulary object detection system, ensuring more efficient training without compromising the model’s effectiveness. To demonstrate the effectiveness of our method, we validate it using the MDETR architecture.
III Method
MDETR is built on the traditional object detection system DETR [20]. DETR is an end-to-end object detection model built with a convolutional residual network backbone and a Transformer Encoder-Decoder [21] architecture. The encoder processes flattened 2D image features from the backbone, while the decoder uses learned object queries, which serve as slots to detect objects in the image. Through cross-attention, the decoder predicts embeddings for each object query, which are then decoded into bounding boxes and class labels. DETR is trained using Hungarian matching to align the predicted objects with ground-truth, utilizing a combination of L1 loss and Generalized IoU [22] for bounding box supervision.
MDETR extends DETR by integrating both visual and textual information into a unified framework. Unlike DETR, which classifies objects into fixed categories, MDETR associates detected objects with spans of text. It encodes images using a ResNet [23] backbone and text via a pre-trained language model (RoBERTa [24]), projecting both into a shared embedding space (ref. Fig. 1). These features are concatenated and processed through a joint transformer encoder. The transformer decoder then cross-attends to this combined representation, predicting object bounding boxes linked to the text.
@@ III Method
For training, MDETR employs two additional key loss functions to align image and text data the soft token prediction loss and the contrastive alignment loss. The soft token prediction loss ($\mathcal{L}_{soft\_token}$) guides the model to predict a uniform distribution over the tokens in the text that correspond to each detected object, rather than predicting discrete class labels.
Figure 1: MDETR Architecture: Visual features are extracted via ResNet and textual features through RoBERTa. Both are projected into a shared embedding space, concatenated, and processed by a transformer encoder-decoder, which predicts object bounding boxes and their alignment with the text.
@@
III-A LightMDETR
III-A1 Architecture
Figure 2: LightMDETR Architecture: Visual and textual features are extracted via frozen ResNet and RoBERTa, then projected into a shared embedding space. A lightweight Universal Projection (UP) encoder, the only trainable component, processes early fused modality features using a learnable ”modality token” $t_{m}$. The UP outputs are concatenated and fed into a Transformer encoder-decoder (DETR) to predict object bounding boxes.
We depict the LightMDETR architecture in Fig. 8. The image is encoded by a frozen ResNet backbone, producing feature vectors $O$, while the text is encoded by a frozen RoBERTa model, yielding feature vectors $T$. Both image and text features are projected into a shared embedding space, with the ”Universal Projection” (UP) module being the only trained component in the backbone. The UP acts as a lightweight encoder, adapting the frozen feature representations for the target task. To handle both modalities, an early fusion method combines the features with a learnable ”modality token” $t_{m}$, specific to each modality (image or text). This approach allows the UP to encode both types of features as follows:
$O_{UP}$ $=UP(O\otimes t_{\text{image}})$ (8)
$T_{UP}$ $=UP(T\otimes t_{\text{text}})$
where $\otimes$ denotes the fusion operation (e.g., addition, multiplication, concatenation, or cross-attention). The outputs $O_{UP}$ and $T_{UP}$ are then concatenated, similar to MDETR, and used as input for the transformer encoder-decoder (DETR) to predict object bounding boxes.
@@ III-A1 Architecture
Figure 3: Architecture of LightMDETR-Plus: LightMDETR-Plus extends LightMDETR 2 by introducing a cross-fusion layer prior to the UP thereby enhancing the model’s representation capabilities.
In MDETR, image and text features are encoded separately and only concatenated before being passed into DETR. However, as shown in [13], early fusion of image and text features can make visual features language-aware, allowing predictions to be conditioned on the text prompt. Building on this idea, we introduce an enhanced version of LightMDETR, called LightMDETR-Plus, shown in Fig. 3. LightMDETR-Plus adds three key components: a cross-fusion layer with Multi-Head Attention (MHA) [21], and two projection layers that refine MHA outputs before they are processed by the UP.
@@
The MHA takes as input the ResNet and RoBERTa encoder outputs, denoted as $O$ and $T$, respectively. The transformations are expressed as:
@@
$T^{(v)}$ $=TW^{(v,T)},$ $O_{F}$ $=\text{SoftMax}(\text{Attn})\cdot T^{(v)}\cdot W^{(\text{out},O)},$
$O^{(v)}$ $=OW^{(v,O)},$ $T_{F}$ $=\text{SoftMax}(\text{Attn}^{\top})\cdot O^{(v)}\cdot W^{(\text{out},T)},$
@@
After applying the cross-fusion mechanism with the Multi-Head Attention approach, a projection is performed using $P_{1}$ and $P_{2}$:
$O_{P_{1}}$ $=P_{1}(O_{\text{F}}+O),$ (10)
$T_{P_{2}}$ $=P_{2}(T_{\text{F}}+T).$
The resulting $O_{P_{1}}$ and $T_{P_{2}}$ are then fed into the UP, following a similar process as in LightMDETR, as described by:
$O_{UP}$ $=UP(O_{P_{1}}\otimes t_{\text{image}}),$ (11)
$T_{UP}$ $=UP(T_{P_{2}}\otimes t_{\text{text}}).$
The proposed lightweight framework for open-vocabulary object detection is modular. Similar to MDETR, we use an end-to-end approach and the same loss function 4 to train both LightMDETR and LightMDETR-Plus. To validate these methods, we compare their performance to MDETR on downstream tasks, including phrase grounding, referring expression comprehension, and segmentation.
@@ IV-A Pre-training
For both LightMDETR and LightMDETR-Plus, we use a frozen pre-trained RoBERTa-base [24] as the text encoder, which consists of 12 transformer layers, each with a 768-dimensional hidden state and 12 attention heads, totaling 125M parameters. The visual backbone is a frozen pre-trained ResNet-101 [23], with 44M parameters. The only trainable component in both models is the UP module (see Fig. 2 and 3), composed of four transformer layers with four attention heads, contributing 4M trainable parameters. In LightMDETR-Plus, projection layers $P_{1}$ and $P_{2}$ add a single transformer layer each, with 787,968 parameters. The modality tokens $t_{image}$ and $t_{text}$ are initialized randomly. By freezing both pre-trained encoders, we reduce the number of trainable backbone parameters from 169M in the original MDETR to 4M in LightMDETR and 5M in LightMDETR-Plus (ref. Fig. 4).
Figure 4: Comparison of trainable backbone parameters (in millions) during training between MDETR, LightMDETR, and LightMDETR-Plus.
For the fusion operation in the UP, as described in Equation 8, we use an addition method. All models are pre-trained for 40 epochs with an effective batch size of 64.
@@ IV-B Dowstream Tasks
We evaluate our method on three downstream tasks: phrase grounding, referring expression comprehension, and segmentation. For a fair comparison, we use the same experimental setup as MDETR. Further details can be found in the original paper.
IV-B1 Phrase grounding
is a task of identifying the fine-grained correspondence between phrases in a sentence and objects (or regions) in an image. We use the Flickr30k entities dataset for this task, and evaluate models performance in terms of Recall@k. For each sentence in the test set, 100 bounding boxes are predicted and use the soft token alignment prediction to rank the boxes according to the score given to the token positions.
Method Val Test
R@1 R@5 R@10 R@1 R@5 R@10
MDETR 82.5 92.9 94.9 83.4 93.5 95.3
LightMDETR 83.98 93.15 94.20 83.87 94.10 95.17
LightMDETR-Plus 84.02 93.56 94.9 83.80 94.66 95.23
TABLE I: Results on the phrase grounding task on Flickr30k dataset.
As shown in Table I, both LightMDETR and its extended version, LightMDETR-Plus, demonstrate competitive performance compared to MDETR. LightMDETR-Plus achieves the highest Recall@1 and Recall@5 on the validation set, with a slight improvement over LightMDETR and MDETR. On the test set, LightMDETR-Plus also outperforms the other models in Recall@5, demonstrating its effectiveness in grounding phrases more accurately. Overall, these results highlight that LightMDETR and LightMDETR-Plus not only reduce the number of trainable parameters but also maintain or slightly improve performance on this task.
Figure 5: An illustration of LightMDETR on modulated detection. The model is designed to identify the root of a phrase as the positive token span, as demonstrated in these figures.
IV-B2 Referring expression comprehension
entails locating an object in an image using a textual description to predict a bounding box. We fine-tune both models on specific datasets—RefCOCO [28], RefCOCO+ [29], and RefCOCOg [30]—for five epochs, while keeping ResNet-101 and RoBERTa frozen. During inference, the models leverage the $\emptyset$ label to rank the 100 predicted bounding boxes, thereby improving the accuracy of object identification based on the provided expression.